OpenStudy (anonymous):

Solve with an integral? ∫ 8^x(1+8^x)^9 dx Should I put a u=1+8^x ? And if I do how do I find the anti-derivative of that and so on?

OpenStudy (anonymous):

to find the anti derivative , write 8 as an exponential with a logarithmn. \[8 =e^\left\{ \ln 8 \right\}\]

OpenStudy (anonymous):

do you recall that \(\frac{d}{dx}b^x=b^x\ln(b)\) ?

OpenStudy (anonymous):

or what @Spacelimbus said, but it might be easier to know that \[\int b^x=\frac{b^x}{\ln(b)}\]

OpenStudy (anonymous):

I was thinking of just making u=1+8^x so du would be 8^x/ln 8 Then somehow go from there?

OpenStudy (anonymous):

You might not need a substitution at all if you apply the formulas that @satellite73 suggested, just multiply out and apply the chain rule to the second term.

OpenStudy (anonymous):

you could always just multiply out

OpenStudy (anonymous):

what @Spacelimbus said no need for a substitution, keep it simple

OpenStudy (anonymous):

Multiply out?

OpenStudy (anonymous):

So like 8^x*1 + (8^2x)^9 ?

OpenStudy (anonymous):

where does the ^9 come from?

OpenStudy (anonymous):

Snap. The problem is: ∫8^x(1+8^x)^9 dx I totally forgot to add that part.

OpenStudy (anonymous):

hehehe, in that case your substitution is a good approach.

OpenStudy (anonymous):

What would my du=? then My u=1+8^x dx du=8^x/ln 8 dx=?

OpenStudy (anonymous):

ok that is wrong i have it backwards

OpenStudy (anonymous):

\[u=1+8^x\] \[du=8^x\ln(8)dx\] \[\frac{du}{\ln(8)}=8^xdx\] and now you are in good shape

OpenStudy (anonymous):

\[\frac{1}{\ln(8)}\int u^9du\] etc

OpenStudy (anonymous):

YES THANK YOU.

OpenStudy (anonymous):

So my final answer I got was 8^x/ln(8) * 9(1+8^x)^8 Correct or no?